OpenStudy (anonymous):

how do you expand 5(3x-1)^2 ?

OpenStudy (mathmagician):

\[=5(9x^2-6x+1)=45x^2-30x+5\]

OpenStudy (anonymous):

square the parens then distribute the 5

OpenStudy (anonymous):

A simple approach if you're just learning algebra is to first expand the brackets: \[5(3x-1){^2}\]\[=5(3x-1)(3x-1)\] Then go along the equation and multiply out each term, remembering to distribute the 5: \[=5\times(3x \times 3x) - 5\times 1\times 3x - 5\times 1 \times 3x + 5\times 1 \times 1\] Now begin to group like terms and simplify: \[=5\times 9x{^2} - 10\times 3x + 5\]\[=45x{^2}-30x+5\]

OpenStudy (anonymous):

First step = 5(9x^2-3x-3x+1) Second step =5(9x^2-6x+1) Third step = 45x^2-30x+5

OpenStudy (anonymous):

thanks !